QUESTION 3

  1. A 54 kg pole vaulter intends to pass over a 6.9-meter high bar. She needs to maintain a minimal speed of 0.79 m/s at the top in order to pass the bar. What is the minimum speed at which she should run in order to attain the height? Neglect air resistance, as well as any energy absorbed
    by the pole.

    Answer in units of m/s and round to one decimal place.

    Hints: Mechanical energy is conserved during vaulting. KE at the bottom is stored in the pole as elastic PE. The elastic PE is turned into the KE+PE at the top where the pole is fully relaxed.
Answer #1

Here,

let the minimum speed needed is u

Using conservation of energy

0.50 * m * v^2 + mgh = 0.50 * m * u^2

0.50 * 0.79^2 + 9.8 * 6.9 = 0.50 * u^2

solving for u

u = 11.7 m/s

the minimum speed needed is 11.7 m/s